OpenStudy (anonymous):

how many positive and negative charges in 250gm of water?

OpenStudy (anonymous):

Water is neutral, but if you can count the electrons (negative charges) and protons (positive charges) as follows: 1 mole of any compound has 6.022*10^23 molecules of that compound. 1 mole of water is 18g.\[250g \left( \frac{ 1mole }{ 18g } \right)\frac{ 6.022*10^{23} molecules}{ mole }=8.364*10^{24}molecules=N\]Each molecule of water has 2 hydrogen molecules, each having 1 proton. Each oxygen atom (of which there's only one in each water molecule) has 8 protons, so the total number of protons is:\[2*1*N+1*8*N=protons\]Since water has no charge the number of electrons equals the number of protons.